The bulk density of the chips was measured by weighing a batch of 195 liters of chips. The measurement result is 42.5kg. The volume was measured to within ± 2.5 liters. At what accuracy should the mass be measured to determine the bulk density to an accuracy of at least ± 3%?
1
Expert's answer
2019-12-13T09:56:41-0500

By the definition, bulk density is


"\rho=\frac{m}{V}"

Let,


"V_0=195\,liters\\[0.3cm]
\Delta V=2.5\,liters\\[0.3cm]
m_0=42.5\,kg"

Then,


"\rho=\overline{\rho}+\Delta\rho,\,where\\[0.3cm]
\overline{\rho}=\frac{m_0}{V_0}"

By the condition of the problem


"\Delta\rho=0.03\cdot\overline{\rho}"

As we know, to find the random error of indirect measurements, you should use the formulas:


"F(x,y,z,\ldots)\longrightarrow\\[0.3cm]
\Delta F=\pm\sqrt{\left(\frac{\partial F}{\partial x}\cdot\Delta x\right)^2+\left(\frac{\partial F}{\partial y}\cdot\Delta y\right)^2+\cdots}"

In our case,


"\rho=\frac{m}{V}\longrightarrow\\[0.3cm]
\Delta\rho=\pm\sqrt{\left(\frac{\partial\rho}{\partial m}\cdot\Delta m\right)^2+\left(\frac{\partial\rho}{\partial V}\cdot\Delta V\right)^2}\longrightarrow\\[0.3cm]
(0.03\cdot\overline{\rho})^2=\left(\frac{\Delta m}{V_0}\right)^2+\left(-\frac{m_0\Delta V}{V_0^2}\right)^2\longrightarrow\\[0.3cm]
\left.\left(\frac{\Delta m}{V_0}\right)^2=\left(\frac{0.03m_0}{V_0}\right)^2-\left(-\frac{m_0\Delta V}{V_0^2}\right)^2\right|\times V_0^2\\[0.3cm]
(\Delta m)^2=(0.03m_0)^2-\left(-\frac{m_0\Delta V}{V_0}\right)^2\longrightarrow\\[0.3cm]
\Delta m=m_0\cdot\sqrt{0.0081-\left(\frac{\Delta V}{V_0}\right)^2}=42.5\cdot\sqrt{0.0081-\left(\frac{2.5}{195}\right)^2}\\[0.3cm]
\Delta m\approx1.15\,(kg)"

ANSWER

"\Delta m\approx 1.15\,kg"